Circular No. 6583

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) K. Birkle, Max-Planck-Institut für Astronomie, Heidelberg; and H.

Böhnhardt, Universitäts Sternwarte, Munich, write: “We have detected ring- and spiral-like coma structures on CN-filtered (388.5 ± 2.3 nm) exposures of C/1995 O1 obtained with the Calar Alto 1.2-m telescope. After adaptive Laplace filtering of the images (via a MIDAS procedure developed by G. Richter, Astronomical Institute, Potsdam), two almost concentric and slightly elliptical rings (axis ratio ∼ 1.1) of diameter ∼ 57′′ and 125′′ along the major axis were found on Feb. 7.245 UT. On Feb. 25.239, another pattern of double rings was observed with diameters 85′′ and 175′′. On both nights, the brightness peak in the coma was displaced to the southwest of the common ring center, and the CN emission appeared weakest in the northern part of the rings. Two strong, straight jets were observed to the south of the central coma peak: one pointed constantly towards p.a. 170^{o}, while the other one changed from p.a. 215^{o} on the first night to p.a. 245^{o} on the second night. On Feb. 25, a shell-like spiral-arm segment extended from the southeast to the northeast coma quadrant at ∼ 130′′ to 250′′ from the central coma peak. The variable jet in the southwest coma quadrant continued in a short spiral arm curved toward the south at ∼ 115′′ from the center. A spiral arclet of similar shape was found in the southwest quadrant at about double the projected distance from the coma center. It is very likely that the ring and spiral arm patterns arise from two different active regions on the nucleus (one for the jet and one for the two rings). Assuming the dust-shell rotation period given by Lecacheux et al. (IAUC 6560), we conclude from the angular separation of the two rings a CN gas-expansion velocity of ∼ 1.2 km/s on Feb. 7 and 1.4 km/s on Feb. 25.”

L. Jorda, Max-Planck Institut für Aeronomie; J. Lecacheux, Observatoire de Paris, Meudon; and F. Colas, Bureau des Longitudes, report: “Further CCD observations of the bright southwestern jet of comet C/1995 O1 at the Pic du Midi 1.05-m telescope confirm the spin period of 11.47 ± 0.05 hr (IAUC 6560) but show a periodic oscillation of this period between 11.20 ± 0.10 hr and 11.65 ± 0.10 hr with a superperiod of 22 ± 2 days. The maximum spin period was reached on ∼ Feb. 14 and Mar. 8, and the minimum near Feb. 27. A complex rotational state of the nucleus is likely to explain this variation, although we are not able to detect the periodic tilt of the spin axis that should be associated with it. A sequence of images covering a whole rotation period is available at URL

http://www.bdl.fr/s2p/comete/halebopp/hbrot.html.”

1997 March 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green